OpenStudy (anonymous):

how do i estimate 14 (square root) to the nearist integer

OpenStudy (anonymous):

It is \[\sqrt{16}\]

OpenStudy (anonymous):

Thanks :D

OpenStudy (anonymous):

find the nearest number to your answer. you want sqrt(14) - the closest whole number would be 16, since sqrt(16) is exactly 4. f(x) = sqrt(x) = x^(1/2) f(16) = sqrt(16) = 4 derive that: f ' (x) = 1/2(x)^(-1/2) = 1/2(sqrt(x)) f ' (16) = 1/2*(sqrt(16)) = 1/2*4 = 1/8 1/8 is the slope of the graph. so. sqrt(14) would be 4 - 2*(1/8) = 4.75 real sqrt(14) is 3.741657 the numbers in the answer represent : 4 is sqrt(16) 1/8 is the slope of the graph 2 is the difference between 16 and 14.

OpenStudy (phi):

you would write down 3*3 = 9 and 3 = sqr(9) 4*4= 16 and 4= sqr(16) then you would say 14 is closer to 16 (2 away) than 9 (5 away) so pick 4

OpenStudy (phi):

in other words \( \sqrt{14} \) will be between \( \sqrt{9}\) and \(\sqrt{16}\) which means \( \sqrt{14} \) will be between 3 and 4 you just have to figure out if you should pick 3 or 4 (if you want the nearest integer)